# Balancing Efficiency and Expressiveness: Subgraph GNNs with Walk-Based Centrality

### v1

@@ 6 Conclusion
We introduced a novel framework, termed HyMN, which combines a subgraph sampling strategy and structural encodings both derived from walk-based centrality measures and, in particular, the Subgraph Centrality by (Estrada & Rodriguez-Velazquez, 2005). We showed that this centrality is a good measure of subgraph importance: for a very limited number of subgraphs it enables competitive performance and outperforms random and learnable selection strategies, approaching full-bag methods. The additional inclusion of centrality-based SEs is also proved to be beneficial both theoretically and in practice, allowing to enhance downstream generalization performance on several real-world benchmarks. Importantly, the strong performance of our method is achieved with a thin computational overhead, making it applicable to a wider spectrum of downstream tasks.
@@
Limitations and Future Work. Our sampling procedure does not take into account already sampled subgraphs unlike methods such as the ones in (Zhao et al., 2022; Bevilacqua et al., 2024). Future work could focus on more general perturbation analyses to give an indication on multi-node marking for higher-order selection policies (Qian et al., 2022) or to quantify the impact of adding subgraphs to a partially populated bag. More sophisticated selection strategies could combine different walk-based centrality measures or consider pairwise structural features.
@@ Acknowledgments
The authors would like to thank Giorgos Bouritsas for insightful early-stage discussions. J.S. is supported by the UKRI CDT in AI for Healthcare http://ai4health.io (EP/S023283/1). M.B. is supported by EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1 and EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1. H.M. is a Robert J. Shillman Fellow and is supported by the Israel Science Foundation through a personal grant (ISF 264/23) and an equipment grant (ISF 532/23). F.F. is funded by the Andrew and Erna Finci Viterbi Post-Doctoral Fellowship. F.F. partly performed this work while visiting the Machine Learning Research Unit at TU Wien led by Prof. Thomas Gärtner.
@@ Appendix A Outline of the Appendix Corpus
  • Appendix C: we report proofs pertaining to Section 4;
  • Appendix D: we provide derivations and additional details on the perturbation analysis of Section 3;
  • Appendix E: we report and discuss further, additional experiments – we compare Subgraph Centrality with other centrality measures and with RWSEs, and we further examine the impact of including CSEs in our model, alongside marking;
  • Appendix F: we comprehensively describe our experimental settings and details, architectural choices and hyper-parameters;
  • Appendix G: we enclose additional timing comparisons.
@@ B.1 The architectural family of Subgraph GNNs
The most popular selection policies are node-based: selected subgraphs are in a bijection with nodes in the original input graph. Prominent policies in this class include node-deletion, ego-networks and node-marking. From an expressiveness perspective, node-marking subsumes the first two policies (Papp & Wattenhofer, 2022; Zhang et al., 2023a). This policy constructs bags of subgraphs as:
@@ B.3 Subsampling bags of subgraphs
To mitigate the aforementioned issue, one possibility is to reduce the number of subgraphs to process, i.e., to lower the impact of $m$ in Equation 7. A convenient approach is to sample a small set $k$ of subgraphs from the bag generated by a predefined policy $\pi$ (Bevilacqua et al., 2021; Zhao et al., 2022; Bevilacqua et al., 2024; Kong et al., 2024; Sun et al., 2021). Essentially, this requires updating Equation 5 as:
@@
Appendix C Claims and proofs
Proposition 1.
@@ Appendix C Claims and proofs
where $\mathbf{W}^{(l)_{1}}$ and $\mathbf{W}^{(l)_{0}}$ are learned weight matrices, and $f^{(l)}$ is an MLP for the $l$-th layer. We refer to the layers in Eqs. 11 and 12 as the MPNN layer and MLP layer, respectively, where the MLP layer is assumed to be a single hidden layer with interleaved ReLU activations.
@@
Using the universal approximation theorem Hornik (1991); Cybenko (1989), $F$ can be approximated to an arbitrary precision using an MLP. Thus, we use the MLP of layer $k+1$ to realize $F$, obtaining the following node matrix,
@@
Theorem 3 (Theorem 1 in Section 4).
@@
Using the notation of Read & Wilson (1998), let $G,G^{\prime}$ be the quartic vertex transitive graphs Qt15 and Qt19 respectively (Here vetrex transitive means that for each pair of nodes there exists a graph automorphism that maps one node to the other, and quartic refers to 4-regular). As these graphs are 4-regular and of the same size, they are 1-WL indistinguishable. In addition, as they are vertex transitive, for each pair of indices $i,j$ we have:
@@
Here the last equalities hold because both graphs have 12 vertices. Thus, to show that $G$ and $G^{\prime}$ are indistinguishable by MPNN + centrality encoding it is enough to show that $\text{trace}(A^{k})=\text{trace}(A^{\prime k})$. As $G$ and $G^{\prime}$ were shown in Brouwer & Spence (2009) to be co-spectral (i.e. their laplacian has the same eigenvalues) and 4-regular, matrices $A$ and $A^{\prime}$ have the same eigenvalues. Thus we have:
@@
Theorem 4 (Theorem 2 in Section 4).
@@
We begin by examining the scenario where $k=1$, meaning that our policy randomly selects subgraphs corresponding to the node with the highest centrality measure. Consider the graph $G$, which is formed by attaching a global node to every vertex of a cyclic graph of length $6$ Next, define $G^{\prime}$ as the graph obtained by attaching a global node to each vertex of two disconnected cyclic graphs, each of length $3$ (The global node is also attached to itself through a self loop). These graphs are displayed in Fig. 5. It can be easily seen that $G$ and $G^{\prime}$ are WL indistinguishable (e.g. by induction). We first prove that both in $G$ and $G^{\prime}$ the global node has the highest centrality measure. This implies that for both graphs, the resulting bag of subgraphs is of size one and is thus equivalent to standard message passing on the graphs (here we can ignore marking as the global nodes have a unique degree and so they can be uniquely identified by standard message passing). This implies that the two graphs are indistinguishable by any Subgraph GNN model $M_{\text{subgraph}}$ which uses a top-$1$ node centrality policy. We then show that the multiset of values of the centrality encoding of each graph is different, showing that it can be distinguished by an MPNN + centrality encoding model. To show that in both graphs the global node has the higher centrality, we first prove the following lemma:
Lemma 1.
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A^{k}_{v,u_{2}}}$ ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}>0}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A^{k+1}_{u_{2},v}=A^{k}_{u_{2},:}\cdot A_{:,v}\geq A^{k}_{u_{2},:}\cdot A_{:,u_{1}}=A^{k+1}_{u_{2},u_{1}}.}$ (17)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A^{k+1}_{u_{2},v}=A_{u_{2},:}\cdot A^{k}_{:,v}\geq A_{u_{2},v}\cdot A^{k}_{v,v}>0.}$ (18)
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}A^{k+1}_{v,v}=A^{k}_{v,u_{1}}\cdot A_{u_{1},v}+\sum_{u\neq u_{1}}A^{k}_{v,u}\cdot A_{u,v}>A^{k}_{u_{1},u_{1}}\cdot A_{u_{1},u_{1}}+\sum_{u\neq u_{1}}A^{k}_{u_{1},u}\cdot A_{u,u_{1}}=A^{k+1}_{u_{1},u_{1}}.}$ (19)
@@
$\tilde{X}^{l-1}=\text{MPNN}^{l;2}\left(\sum_{j=1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n}}A^{j},\sum_{j=1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}n}}X^{l-1;j}\right),$ (23)
@@
Appendix D On marking-induced perturbations and the number of walks
@@
We can readily apply Proposition 3 and leverage the fact that marking only induces a unit-norm feature perturbation to get:
@@
The above bound is easily extended to consider a closely related analysis: the impact of adding a single node-marked subgraph $S$ to a bag formed by the original graph only. This analysis would enlighten us on the impact a single subgraph addition in the case of “augmented policies”, which we use throughout our experiments in Section 5, see Section F.1.
@@
where $(\mathcal{B})$ can be upper-bounded, again, by Proposition 3101010We have allowed a little abuse of notation here by using $S$ to refer to the “subgraph” obtained by marking node $S$ in $G$..
@@
Appendix E Additional Experiments
E.1 Comparison to other centrality measures

### v2

@@ 6 Conclusions
We introduced a novel Subgraph GNN method, termed HyMN, which strikes a compelling balance between efficiency and expressiveness by combining a subgraph sampling strategy and structural encodings both derived from walk-based centrality measures. HyMN attains strong empirical performance with an extremely thin computational overhead, making it applicable to a wide spectrum of downstream tasks featuring large graphs, even tens or hundreds of times larger than those full-bag Subgraph GNNs can possibly process. In our paper, we showed, in particular, that the Subgraph Centrality by Estrada & Rodriguez-Velazquez (2005) is a good measure of subgraph importance: for a very limited number of subgraphs it enables competitive performance and outperforms random and learnable selection strategies. We also proved that the additional inclusion of centrality-based SEs is beneficial both theoretically and in practice, allowing to enhance discriminative power and downstream generalization performance on several real-world benchmarks.
@@
Limitations and future work. Our sampling procedure does not take into account already sampled subgraphs unlike methods such as the ones in (Zhao et al., 2022; Bevilacqua et al., 2024). Future work could focus on more general perturbation analyses to give an indication on multi-node marking for higher-order selection policies (Qian et al., 2022) or to quantify (i) the optimality of the designed sampling strategy, (ii) the impact of adding subgraphs to a partially populated bag. More sophisticated selection strategies could combine different walk-based centrality measures or consider pairwise structural features.
@@ Acknowledgments
The authors would like to thank Giorgos Bouritsas for insightful early-stage discussions and Moshe Eliasof for late stage experimental insights on PolicyLearn. J.S. is supported by the UKRI CDT in AI for Healthcare http://ai4health.io (EP/S023283/1). Y.E. is supported by the Zeff PhD Fellowship; G.B.S. by the Jacobs Qualcomm PhD Fellowship. M.B. is supported by EPSRC Turing AI World-Leading Research Fellowship No. EP/X040062/1 and EPSRC AI Hub on Mathematical Foundations of Intelligence: An “Erlangen Programme” for AI No. EP/Y028872/1. H.M. is a Robert J. Shillman Fellow and is supported by the Israel Science Foundation through a personal grant (ISF 264/23) and an equipment grant (ISF 532/23). F.F. conducted this work supported by an Andrew and Erna Finci Viterbi and, partly, by an Aly Kaufman Post-Doctoral Fellowship. F.F. partly performed this work while visiting the Machine Learning Research Unit at TU Wien led by Prof. Thomas Gärtner.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Outline of the Appendix Corpus
  • Appendix C: we report a full algorithmic description of our HyMN approach;
  • Appendix D: we report proofs pertaining to Section 4;
  • Appendix E: we provide derivations and additional details on the perturbation analysis of Section 3;
  • Appendix F: we report and discuss further, additional experiments – we additionally run HyMN on the ZINC dataset, we explore more backbone MPNN architectures, we compare Subgraph Centrality with other centrality measures and with RWSEs, and we further examine the impact of including CSEs in our model, alongside marking;
  • Appendix G: we comprehensively describe our experimental settings and details, architectural choices and hyper-parameters;
  • Appendix H: we enclose additional timing comparisons.
@@ B.1 The Architectural Family of Subgraph GNNs
The most popular selection policies are node-based: selected subgraphs are in a bijection with nodes in the original input graph. Prominent policies in this class include node-deletion, ego-networks and node-marking. From an expressiveness perspective, node-marking subsumes the first two policies (Papp & Wattenhofer, 2022; Zhang et al., 2023a), other than uncovering a connection between node-based Subgraph GNNs and node-individualization algorithms for graph isomorphism testing (Dupty & Lee, 2022). Node-marking constructs bags of subgraphs as:
@@ B.2 Computational Complexity Aspects
Figure 4: Depiction of a Subgraph GNN endowed with a subgraph sampling strategy.
@@ B.3 Subsampling Bags of Subgraphs
To mitigate the aforementioned issue, one possibility is to reduce the number of subgraphs to process, i.e., to lower the impact of $m$ in Equation 7. A convenient approach is to sample a small set $k$ of subgraphs from the bag generated by a predefined policy $\pi$ (Bevilacqua et al., 2021; Zhao et al., 2022; Bevilacqua et al., 2024; Kong et al., 2024; Sun et al., 2021) (see Figure 4). Essentially, this requires updating Equation 5 as:
@@
Appendix C Hybrid Marking Networks in Algorithmic Form
Algorithm 1 Hybrid Marking Network
0:  Graph $G=(A,X)$, Subgraph GNN $f$, Max walk-length $K$, Number of marks $T$
1:  $C^{\mathsf{CSE}}_{v,k}\leftarrow 0\quad\forall v\in G,k\in[K]$triangleright SE init.
2:  $M_{v,t}\leftarrow 0\quad\forall v\in G,t\in[T]$triangleright Mark init.
3:  $B\leftarrow I$
4:  for $k\in[K]$ do
5:     for $v\in G$ do
6:        $C^{\mathsf{CSE}}_{v,k}\leftarrow\nicefrac{{B_{vv}}}{{k!}}$triangleright Compute SEs
7:     end for
8:     $B\leftarrow B\cdot A$
9:  end for
10:  $\tilde{C}^{SC}\leftarrow\sum_{k}C^{\mathsf{CSE}}_{:,k}$triangleright Estimate SC
11:  $\mathcal{M}\leftarrow\text{select-top}(\tilde{C}^{SC},T)$triangleright Select nodes
12:  for $t\in[T]$ do
13:     $M_{\mathcal{M}[t],t}\leftarrow 1$triangleright Mark nodes
14:  end for
15:  $y_{G}\leftarrow f(A,X\oplus C^{\mathsf{CSE}},M)$triangleright Forward-pass
16:  return $y_{G}$
Algorithm 1 describes our HyMN approach, i.e., how graph representations are obtained starting from an input graph $G=(A,X)$, a backbone Subgraph GNN $f$, and hyper-parameters $K,T$ [^7] Note that, as reported in Algorithm 1, we perform a select-top operation (line 11). In some cases, which we found to rarely occur in practice, depending on the value of $T$, this operation has to arbitrarily break ties between nodes with the same exact centrality values. This tie breaking procedure can additionally be either side-stepped by marking all nodes with the same centrality value or be made equivariant by accounting for the ordering between colours from an auxiliary DS-WL-like refinement (Bevilacqua et al., 2021)., specifying, respectively, the maximum walk-length and the number of marks/subgraphs to select.
Appendix D Claims and Proofs
Proposition D.1.
@@ Appendix D Claims and Proofs
where $\mathbf{W}^{(l)_{1}}$ and $\mathbf{W}^{(l)_{0}}$ are learned weight matrices, and $f^{(l)}$ is an MLP for the $l$-th layer. We refer to the layers in Equations 11 and 12 as the MPNN layer and MLP layer, respectively, where the MLP layer is assumed to be a single hidden layer with interleaved ReLU activations.
@@
Using the universal approximation theorem (Hornik, 1991; Cybenko, 1989), $F$ can be approximated to an arbitrary precision using an MLP. Thus, we use the MLP of layer $k+1$ to realize $F$, obtaining the following node matrix,
@@
Theorem D.2 (Theorem 4.1 in Section 4).
@@
Using the notation of (Read & Wilson, 1998), let $G,G^{\prime}$ be the quartic vertex transitive graphs Qt15 and Qt19 respectively (Here vetrex transitive means that for each pair of nodes there exists a graph automorphism that maps one node to the other, and quartic refers to 4-regular). As these graphs are 4-regular and of the same size, they are 1-WL indistinguishable. In addition, as they are vertex transitive, for each pair of indices $i,j$ we have:
@@
Here the last equalities hold because both graphs have 12 vertices. Thus, to show that $G$ and $G^{\prime}$ are indistinguishable by MPNN + centrality encoding it is enough to show that $\text{trace}(A^{k})=\text{trace}(A^{\prime k})$. As $G$ and $G^{\prime}$ were shown in (Brouwer & Spence, 2009) to be co-spectral (i.e. their laplacian has the same eigenvalues) and 4-regular, matrices $A$ and $A^{\prime}$ have the same eigenvalues. Thus we have:
@@
Theorem D.3 (Theorem 4.2 in Section 4).
@@
We begin by examining the scenario where $k=1$, meaning that our policy randomly selects subgraphs corresponding to the node with the highest centrality measure. Consider the graph $G$, which is formed by attaching a global node to every vertex of a cyclic graph of length $6$ Next, define $G^{\prime}$ as the graph obtained by attaching a global node to each vertex of two disconnected cyclic graphs, each of length $3$ (The global node is also attached to itself through a self loop). These graphs are displayed in Figure 6. It can be easily seen that $G$ and $G^{\prime}$ are WL indistinguishable (e.g. by induction). We first prove that both in $G$ and $G^{\prime}$ the global node has the highest centrality measure. This implies that for both graphs, the resulting bag of subgraphs is of size one and is thus equivalent to standard message passing on the graphs (here we can ignore marking as the global nodes have a unique degree and so they can be uniquely identified by standard message passing). This implies that the two graphs are indistinguishable by any Subgraph GNN model $M_{\text{subgraph}}$ which uses a top-$1$ node centrality policy. We then show that the multiset of values of the centrality encoding of each graph is different, showing that it can be distinguished by an MPNN + centrality encoding model. To show that in both graphs the global node has the higher centrality, we first prove the following lemma:
Lemma D.4.
@@
$A^{k}_{v,u_{2}}$ $>0$
@@
$A^{k+1}_{u_{2},v}=A^{k}_{u_{2},:}\cdot A_{:,v}\geq A^{k}_{u_{2},:}\cdot A_{:,u_{1}}=A^{k+1}_{u_{2},u_{1}}.$ (17)
@@
$A^{k+1}_{u_{2},v}=A_{u_{2},:}\cdot A^{k}_{:,v}\geq A_{u_{2},v}\cdot A^{k}_{v,v}>0.$ (18)
@@
$A^{k+1}_{v,v}=A^{k}_{v,u_{1}}\cdot A_{u_{1},v}+\sum_{u\neq u_{1}}A^{k}_{v,u}\cdot A_{u,v}>A^{k}_{u_{1},u_{1}}\cdot A_{u_{1},u_{1}}+\sum_{u\neq u_{1}}A^{k}_{u_{1},u}\cdot A_{u,u_{1}}=A^{k+1}_{u_{1},u_{1}}.$ (19)
@@
$\tilde{X}^{l-1}=\text{MPNN}^{l;2}\left(\sum_{j=1}^{n}A^{j},\sum_{j=1}^{n}X^{l-1;j}\right),$ (23)
@@
Appendix E On Marking-Induced Perturbations and the Number of Walks
@@
We can readily apply Proposition E.1 and leverage the fact that marking only induces a unit-norm feature perturbation to get:
@@
The above bound is easily extended to consider a closely related analysis: the impact of adding a single node-marked subgraph $S$ to a bag formed by the original graph only. This analysis would enlighten us on the impact a single subgraph addition in the case of “augmented policies”, which we use throughout our experiments in Section 5, see Section G.1.
@@
where $(\mathcal{B})$ can be upper-bounded, again, by Proposition E.1111111We have allowed a little abuse of notation here by using $S$ to refer to the “subgraph” obtained by marking node $S$ in $G$..
@@
E.1 Perturbation Analysis on an Additional Dataset
As outlined in Section 3.2, we explore the amount of perturbation from different marking schemes on the output of an untrained GIN model. We have shown in Figure 1, the effect of these marking schemes on the MUTAG dataset. Here, in Figure 7(a) and in Figure 7(b), we additionally show the results for the NCI1 dataset and for Erdös-Renyi (ER) graphs. The ER graphs are consistent with our correlation analysis experiment Section 3.2, where we randomly generate $100$ ER graphs, each with $N=20$ nodes and wiring probability $p=0.3$.
(a) NCI1 dataset alterations.
(b) Erdös-Renyi graph alterations.
Figure 7: Plots showing the alteration of graph representations using GIN by adding an additional node-marked subgraph with (i) the highest centrality, (ii) the lowest centrality, and (iii) a random marking for both the NCI1 dataset (left) and Erdös-Renyi graphs (right).
Appendix F Additional Experiments
F.1 Additional Experiments with Other Backbones
We have already experimented with GCN as the backbone on the Peptides datasets. Additionally, we run GCN as a backbone on OGB molecular datasets (see Table 6). Interestingly, we find that a GCN backbone achieves the best results on MOLBACE. Moreover, we perform experiments on yet another backbone: GatedGCN (Bresson & Laurent, 2017). This complements backbones used in the Graph Transformer literature such as GPS. The results on peptides and MolHIV are shown in Table 7. These results further show that the walk-based centrality structural encoding can improve performance over the backbone MPNN, and we can improve performance even further by using it for sampling subgraphs.
Table 6: Results on OGB benchmarks using GCN as the backbone architecture.
Model MolHIV MolBace MolTox
GCN 76.06 ±$\pm 0.97$ 79.15 ±$\pm 1.44$ 75.29 ±$\pm 0.69$
HyMN (GCN, T=2) 76.82 ±$\pm 1.20$ 83.21 ±$\pm 0.51$ 76.07 ±$\pm 0.50$
Table 7: Results on the Peptides and MolHIV benchmarks using GatedGCN as the backbone architecture.
Model MolHIV Peptides-Func Peptides-Struct
GatedGCN 78.27 ±$\pm 0.65$ 0.6558 ±$\pm 0.0068$ 0.2497 ±$\pm 0.0007$
GatedGCN + CSE 80.33 ±$\pm 0.56$ 0.6611 ±$\pm 0.0058$ 0.2482 ±$\pm 0.0010$
HyMN (GatedGCN, T=2) w/out CSE 79.02 ±$\pm 0.91$ 0.6723 ±$\pm 0.0079$ 0.2473 ±$\pm 0.0013$
HyMN (GatedGCN, T=2) 81.07 ±$\pm 0.21$ 0.6788 ±$\pm 0.0052$ 0.2471 ±$\pm 0.0007$
F.2 Results on ZINC and Additional Model Comparisons on MolHIV
We experimented with the ZINC-12K molecular dataset (Sterling & Irwin, 2015; Gómez-Bombarelli et al., 2018), where we maintain a 500k parameter budget, in line with previous works.
Table 8: Test results on ZINC. The first and second best results are color-coded.
Method ZINC ($\downarrow$)
GCN 0.321 ±$\pm 0.009$
GIN 0.163 ±$\pm 0.004$
GSN 0.101 ±$\pm 0.010$
CIN 0.079 ±$\pm 0.006$
GPS 0.070 ±$\pm 0.004$
GINE-MLP-Mixer 0.073 ±$\pm 0.001$
GINE-ViT 0.085 ±$\pm 0.004$
FULL 0.087 ±$\pm 0.003$
OSAN 0.177 ±$\pm 0.016$
POLICY-LEARN (T = 2) 0.120 ±$\pm 0.003$
RANDOM (T = 2) 0.136 ±$\pm 0.005$
GIN+CSE 0.092 ±$\pm 0.002$
HyMN (GIN, T=1) 0.080 ±$\pm 0.003$
HyMN (GIN, T=2) 0.083 ±$\pm 0.002$
Table 9: Test results on MOLHIV. The first and second best results are color-coded.
Method MOLHIV ($\uparrow$)
GCN 76.06 ±$\pm 0.97$
GIN 75.58 ±$\pm 1.40$
Reconstr. GNN 76.32 ±$\pm 1.40$
DSS-GNN 76.78 ±$\pm 1.66$
GSN 80.39 ±$\pm 0.90$
CIN 80.94 ±$\pm 0.57$
GPS 78.80 ±$\pm 1.01$
GINE-MLP-Mixer 79.97 ±$\pm 1.02$
GINE-ViT 77.92 ±$\pm 1.42$
Nested-GNN 78.34 ±$\pm 1.86$
GNN-AK+ 79.61 ±$\pm 1.19$
SUN 80.03 ±$\pm 0.55$
FULL 76.54 ±$\pm 1.37$
POLICY-LEARN (T = 2) 79.13 ±$\pm 0.60$
RANDOM (T = 2) 77.55 ±$\pm 1.24$
GIN+CSE 77.44 ±$\pm 1.87$
HyMN (GIN, T=1) 80.36 ±$\pm 1.23$
HyMN (GIN, T=2) 81.01 ±$\pm 1.17$
We compared the test MAE using HyMN to other sampling approaches (Qian et al., 2022; Bevilacqua et al., 2024), a full-bag Subgraph GNN, a Graph Transformer (GPS) (Rampášek et al., 2022), two expressive GNN baselines (GSN, CIN) (Bouritsas et al., 2022; Bodnar et al., 2021a) and the two alternative architectures Graph-ViT and Graph MLP-Mixers by (He et al., 2023). Results are presented in Table 8. We additionally report results on MOLHIV in Table 9, often used as a reference test-bed for most full-bag methods (Bevilacqua et al., 2021; Frasca et al., 2022; Zhang & Li, 2021; Cotta et al., 2021; Zhao et al., 2022). We observe that our hybrid method can outperform full-bag Subgraph GNNs as well as previously proposed subsampling based approaches. Additionally, we perform competitively with CIN, which takes into account higher-order interactions and explicitly models ring-like structures. Non-message-passing based architectures can perform slightly better in absolute terms, but for higher run-time (see comparison with GPS in Appendix H). We highlight that additionally using subgraphs can outperform purely using the centrality-based encodings (Q4). In particular, the two approaches in tandem deliver exceptional performance on MolHIV, for $T=2$.
F.3 Minimum Centrality Marking for Counting Substructures
As highlighted in Section 3.2, low-centrality nodes perturb representations only to a limited extent, causing the method to work in a regime close to a vanilla $1$-WL GNN. We also empirically observe that the perturbations induced by marking high-centrality nodes correlate much more with potentially predictive features, such as subgraph counts. In order to further evidence this point, we have run experiments for counting triangle and $4$-cycle substructures, studying the impact of sampling marked subgraphs associated with the minimum values of the Subgraph Centrality (SC). We report the results in Table 10 and Table 11, along with those of random sampling and max-centrality sampling (MAE, the lower the better). We observe that min-centrality sampling is indeed outperformed by the two other strategies on all settings, further suggesting the limitations of marking nodes with low subgraph centrality.
Table 10: Performance comparison for different policies on triangle subgraphs.
Policy 10 Subgraphs 20 Subgraphs 30 Subgraphs
Min SC 0.78 0.52 0.43
Random 0.62 0.48 0.40
Max SC 0.20 0.10 0.03
Table 11: Performance comparison for different policies on 4-cycle subgraphs.
Policy 10 Subgraphs 20 Subgraphs 30 Subgraphs
Min SC 0.74 0.63 0.41
Random 0.59 0.45 0.36
Max SC 0.38 0.12 0.08
F.4 Comparison to Other Centrality Measures